OpenStudy (anonymous):

Find the slope m of the line that passes through the points (–3, 1) and (5, 6). slope m = –5/8 slope m = 5/8 slope m = 8/5 slope m = –8/5

OpenStudy (anonymous):

you use the slope formula m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

so it would be 5/8

OpenStudy (anonymous):

thanks again :)

OpenStudy (anonymous):

6 - 1 = 5 5 - (-3) = 8 so, yeah... the slope is 5/8.